Present: Rich Bennett, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Everett Phillips, Hilton Soto, Ying TangAnnouncements
Review Minutes of 8/16/99
- Word from the stadium is that there is one applicant for the Network Coordinator position. The resume is enroute.
- The new Systems Analyst, Hilton Soto, started working today.
Ongoing Systems Updates
- Nets #4 - LeiLani made another pass through LAC - it is still a configuration problem. Machines need to be brought down & software reinstalled, but they don't want to do that during the first week of classes.
- Office 2000 should be arriving soon. It will be distributed from the server only. We'll have 300 licenses for it, which should cover all of the staff.
- The Office 2000 order does NOT include FrontPage, contrary to earlier indication. FrontPage will be individually licensed - 55 copies are on order, to be distributed as needed.
- HD supply: about 20 are remaining - should be fine, but Everett will review.
- Training session has been completed to mixed reviews - there seems to be a division of opinion between the librarians & USPS. The instructor was well liked and very knowledgeable. The biggest complaint seemed to be that there was not enough emphasis on Windows'98 and more hands-on training on the hardware side would have helped. The post-training test was a take-home, closed book test.
- The next liaison meeting will be Sept. 2. Topics will include feedback on the training and opinions about whether to continue the liaison meetings or not.
- LMG will need to review the feedback on the training session to determine the cost effectiveness.
Software Projects Status
- CSE reconfiguration - occurred last week - added RSM, took CIRCA & CIS department off - this will reduce the load and provide more horse power to the main router. MSL & West are still hooked in, we hope to have West moved over to the Criser router soon.
- NERDC dial-up access 846 phone number is unstable right now. It is recommended that you switch over to the new number immediately.
- The wireless tap is to be put into Special Collections today. This is an experimental installation - we're going to try it out there for a week, then move it to another location (MSL & LAD). The new campus telephone contract may include the wireless technology. It will probably take 4-6 weeks to determine how we'll approach this situation.
- IR database revisions - Bobbie says it's working well but she still would like Ying to make a few more changes.
- Ying has been working on an OPS Task Tracker for Resource Services & Preservation. It may eventually be distributed library-wide. It is based on Personnel records by timecards. Bill notes to be sure the database includes stadium OPS. Since Evelyn is the only one who inputs the actual updates to the records, it would be a good idea to make sure that someone else in Personnel is knowledgeable about it as well.
- Mellon 2 - Winston set David up with an Access database - but they are apparently still waiting for funding.
- Richard Phillips is experiencing a resolution problem with his screen when using Mellon 1. Winston added a T-sizer, but it fought with the TIFF viewer. A newer version of the T-sizer should resolve this problem. We'll order the new version today.
- Hilton is coming up to speed on Delphi - he'll probably start with Business Services dbf files.
- Borland has been polling developers for tools they would like to have & Linux users are requesting Delphi. Delphi version 5 is out.
- Lindsay Penrose will be officially taking Scott's position (ordering) as of Monday, Aug. 30.
- Front Desk students for the Fall term are: Runa, Becky Leidy (returning) & Kelly (new hire - former student, Bobby Costa's sister).
- Leo Petrone will no longer be working with us.
- Suzy Shaw will be out all this week, back on Aug. 30.
- Bill will be out during the afternoon of Wed. Aug 25.
Back to the Systems Homepage
Last updated February 18, 2000
by Debra Harris